a dairy farmer has two types of milk. How many gallons of a 2% milk fat solution and how many gallons of a 4.5% milk fat solution should be mixed to obtain 500 gallons of a 2.9% milk fat solution

Answer by checkley79(3341)

.045X+.02(500-X)=.029*500
.045X+10-.02X=14.5
.025X=14.5-10
.025X=4.5
X=4.5/.025
X=180 GALLONS OF 4.5% IS USED.
500-180=320 GALLONS OF 2% IS USED.
PROOF:
.045*180+.02*320=.029*500
8.1+6.4=14.5
14.5=14.5